# A Structured Neural ODE Approach for Real Time Evaluation of AC Losses in 3D Superconducting Tapes

### v1

@@ 5.2 Training
Once the approximated operator $\mathbf{R}_{\psi}$ has been defined, the next step is to train it. A naive approach would be to collect evaluations of $\mathbf{R}$ for different values of $\mathbf{i}$ and then use them to solve a regression problem. However, this strategy is (i) intrusive, since it requires access to snapshots of the nonlinearity, which is often infeasible when dealing with commercial solvers, and (ii) less effective than training on entire trajectories. In fact, as demonstrated in [42], trajectory-based training is superior to local derivative fitting in terms of robustness and generalization capabilities.
For this reason, we adopt a discretize-then-optimize approach, where the time discretization is embedded into the learning process and the parameters are updated by backpropagation through the solver. The procedure to advance the reduced system from $t_{n}$ to $t_{n+1}$ is as follows:
@@
This training procedure follows the Neural ODEs [34] paradigm, where the dynamics are unrolled across the temporal domain and the parameters are optimized end-to-end. The loss is defined as the Mean Squared Error (MSE) by the predicted states and the true states obtained by simulating the FOM. Figure 2 provides a graphical interpretation of this training process.
@@ 6.1 Data Generation
The resulting model features $n_{e}=4518$ edges and $n_{f}=3140$ faces, yielding a total of 7658 Degrees of Freedom (DOFs). To generate the dataset used for POD-DEIM reduction and for training the structured Neural ODE, five transient simulations were performed. Each transient corresponds to a sinusoidal excitation at 50 Hz with different amplitudes. This setup reflects the physical scenario under consideration, where the external magnetic field induced by surrounding cables is sinusoidal at 50 Hz, and its amplitude varies with both current magnitude and cable distance. To evaluate the generalization capability of the proposed approach, additional transient simulations were conducted with varying amplitudes—both within and outside the training distribution—as well as with different excitation frequencies. The corresponding excitation parameters defining the training and testing datasets are reported in Table 1.
@@ 60 Hz 20 mT
Table 1: Overview of the excitation parameters adopted for dataset generation. The table reports the combinations of frequencies and amplitudes used in the training simulations, as well as those employed for testing within and outside the training distribution.
@@ 6.2 POD-DEIM Practical Implementation
In the implementation of the proposed reduction framework, the training snapshots were used for constructing the snapshot matrices. This procedure yielded a total of $n_{s}=2800$ snapshots for the current, potential, and nonlinear term. For the current field, the dominant modes were first selected based on the energy retained, resulting in $r_{i}=48$ retained modes. For the potential field, only the energy retention criterion was applied, leading to the selection of the first five modes, i.e., $r_{\varphi}=5$. The singular values decay for the two variables are shown in Figure 3.
@@ 6.3 Structured Neural ODE Practical Implementation
The neural network $\mathbf{R}_{\psi}$ is implemented as a fully connected feed-forward architecture with $L_{h}=4$ hidden layers, each consisting of 140 neurons. As a preprocessing step, the absolute value of the reduced current $|\mathbf{i}_{r}|$ is fed to the network, reflecting the fact that resistivity—and therefore the $\mathbf{R}_{\psi}$—depends on the magnitude of the current rather than its sign. All the layers employ selu [43] activation functions. The training dataset considers the very same transients used for constructing the POD-DEIM snapshot matrix. This dataset is split into fixed-length sequences that are organized into mini-batches, and the model is trained using the Adam optimizer [44].
@@ 6.4 Accuracy
The accuracy of each reduced-order model was assessed by comparing the predicted current density with the reference solution obtained from the FOM. Specifically, the instantaneous absolute error was computed as the absolute difference between the ROM-predicted current and the corresponding FOM value. In the time-dependent plots, the mean, 95th-percentile, and maximum errors were evaluated across all current states at each time instant, thereby illustrating the temporal evolution of the error. Conversely, in the summary tables, the reported mean, 95th-percentile, and maximum values were computed over all current states and all time instants of the transient simulations.
@@ 6.4.2 Outside-the-distribution Validation
To evaluate the performance of the ROM s outside the training distribution, two additional transient simulations were performed at 50 Hz with excitation amplitudes of 10 mT and 30 mT—values lying below and above the training range (13–24 mT), respectively. The corresponding quantitative results, summarized in Table 3, indicate that the accuracy of the structured Neural ODE decreases noticeably, whereas the POD-DEIM ROM maintains performance comparable to the within-distribution case.
@@ 6.4.3 Different Frequency Validation
To further assess the generalization capability of the proposed approach, additional simulations were carried out at an excitation amplitude of 20 mT—within the training range—but with frequencies of 40, 60, and 100 Hz, differing from the 50 Hz training frequency. The quantitative results, summarized in Table 4, demonstrate that the structured Neural ODE achieves very good accuracy across all tested frequencies, with errors of the same order of magnitude as in the 50 Hz validation and consistently lower than those of the POD-DEIM ROM.
@@ 6.5 Computational Cost
To compare the computational cost of the two ROM s, we evaluate the number of Floating-Point Operations (FLOPs) required to advance the system by one time step. The time-stepping scheme used to simulate the POD-DEIM ROM employs Newton iterations at each step, with residual and Jacobian evaluations at each iteration—an approach necessitated by the strong nonlinearity of the problem. In both cases, an optimal implementation scenario is assumed.
@@ 6.5.2 Structure Neural ODE ROM Computational Cost
The computational cost for advancing the state in the POD-DEIM reduced-order model includes evaluating the neural network, assembling the left- and right-hand sides through matrix-vector operations, and solving the resulting linear system.
@@ 6.5.3 Results of the Evaluations
For a single time step, the POD-DEIM ROM requires approximately $4.5\times 10^{7}$ operations, while the structured Neural ODE ROM requires only $2.6\times 10^{5}$, corresponding to a theoretical speed-up of $\times 176$. However, achieving this theoretical performance depends on highly optimized implementations. While optimizing the Neural ODE implementation is relatively straightforward—requiring only efficient neural network evaluation and basic algebraic operations—achieving an optimal DEIM implementation is more challenging. In fact, when using DEIM, the FOM assembly routines are still called on the fly during the reduced-order simulation. To reach optimal efficiency, these routines themselves would need to be carefully optimized, which is a nontrivial task in practice. In our implementation, the practical speed-up reaches $\times 1500$, reducing computation from 5 minutes to just 0.2 seconds for a full transient reconstruction. These results are summarized in Table 5.
@@ 6.7 Post-processing: Power losses
As a post-processing step, we evaluate the AC losses in the CORC® cable, which serve as input for a thermal model aimed at predicting the cable’s temperature distribution—a key quantity for quench prediction in both design and control. The AC losses are computed as:
@@ 7 Discussion
Overall, these considerations suggest that while POD-DEIM may be advantageous in simpler or data-scarce scenarios, the structured Neural ODE offers superior speed, flexibility, and precision in settings where sufficient data are available and intrusiveness must be minimized.

### v2

@@ 5.2 Training
Once the approximated operator $\mathbf{R}_{\psi}$ has been defined, the next step is to train it. A naive approach would be to treat the problem as a standard regression task: one would simulate the FOM, project the resulting current snapshots onto the POD basis to obtain reduced states $\big\{\mathbf{i}_{r}^{(n)}\big\}$, evaluate (and, if needed, project) the corresponding nonlinear operators to form the reduced resistance matrices $\big\{\mathbf{R}_{r}(\mathbf{i}_{r}^{(n)})\big\}$, and then train the network by pointwise regression of the mapping $\mathbf{i}_{r}\mapsto\mathbf{R}_{r}(\mathbf{i}_{r})$. However, this strategy is (i) intrusive, since it requires access to snapshots of the nonlinearity, which is often infeasible when dealing with commercial solvers, and (ii) less effective than training on entire trajectories. In fact, as demonstrated in [34], trajectory-based training is superior to local derivative fitting in terms of robustness and generalization capabilities.
For this reason, we adopt a discretize-then-optimize approach, where the time discretization is embedded into the learning process, and the parameters are updated by backpropagation through the solver. The procedure to advance the reduced system from $t_{n}$ to $t_{n+1}$ is as follows:
@@
This training procedure follows the Neural ODE [11] paradigm, where the dynamics are unrolled across the temporal domain and the parameters are optimized end-to-end. The loss is defined as the Mean Squared Error (MSE) by the predicted states and the true states obtained by simulating the FOM. Figure 2 provides a graphical interpretation of this training process.
@@ 6.1 Data Generation
The resulting model features $n_{e}=4518$ edges and $n_{f}=3140$ faces, yielding a total of 7658 Degrees of Freedom (DOFs). To generate the dataset used for POD-DEIM reduction and for training the Structured Neural ODE, five transient simulations were performed. Each transient corresponds to a sinusoidal excitation at 50 Hz with different amplitudes. This setup reflects the physical scenario under consideration, where the external magnetic field induced by surrounding cables is sinusoidal at 50 Hz, and its amplitude varies with both current magnitude and cable distance. To evaluate the generalization capability of the proposed approach, additional transient simulations were conducted with varying amplitudes–both within and outside the training distribution–as well as with different excitation frequencies. The corresponding excitation parameters defining the training and testing datasets are reported in Table 1.
@@ 60 Hz 20 mT
Table 1: Overview of the excitation parameters adopted for dataset generation.
@@ 6.2 POD-DEIM Practical Implementation
In the implementation of the proposed reduction framework, the training snapshots were used for constructing the snapshot matrices. This procedure yielded a total of $n_{s}=2800$ snapshots for the current, potential, and nonlinear term. For the current field, the dominant modes were selected based on the energy retained, resulting in $r_{i}=48$ retained modes. Also for the potential field, the energy retention criterion was applied, leading to the selection of the first five modes, i.e., $r_{\varphi}=5$. The singular values decay for the two variables are shown in Figure 3.
@@ 6.3 Structured Neural ODE Practical Implementation
The neural network $\mathbf{R}_{\psi}$ is implemented as a fully connected feed-forward architecture with $L_{h}=4$ hidden layers, each consisting of 140 neurons. As a preprocessing step, the absolute value of the reduced current $|\mathbf{i}_{r}|$ is fed to the network, reflecting the fact that resistivity–and therefore the $\mathbf{R}_{\psi}$–depends on the magnitude of the current rather than its sign. All the layers employ selu [22] activation functions. The training dataset considers the very same transients used for constructing the POD-DEIM snapshot matrix. This dataset is split into fixed-length sequences that are organized into mini-batches, and the model is trained using the Adam optimizer [21].
@@ 6.4 Accuracy
The accuracy of each reduced-order model was assessed by comparing the predicted current density with the reference solution obtained from the FOM. Specifically, the instantaneous absolute error was computed as the absolute difference between the ROM-predicted current and the corresponding FOM value. In the time-dependent plots, the mean, 95th-percentile, and maximum errors were evaluated across all current states at each time instant, thereby illustrating the temporal evolution of the error. Conversely, in the summary tables, the reported mean, 95th-percentile, and maximum values were computed over all current states and all time instants of the transient simulations. Moreover, the last row of the accuracy tables reports the relative discrepancy between the predicted and reference current trajectories, measured through the Frobenius norm:
$\text{err}_{\lVert\cdot\rVert_{F}}=\frac{\lVert\mathbf{I}-\mathbf{I}_{\text{true}}\rVert_{F}}{\lVert\mathbf{I}_{\text{true}}\rVert_{F}},$ (30)
where $\mathbf{I}$ denotes the matrix collecting the full-order current states at all time instants reconstructed from the ROM solution (i.e., obtained by projecting the reduced currents back to the full space), and $\mathbf{I}_{true}$ is defined analogously using the FOM solution. This quantity provides a global measure of relative error over the entire transient. We note that a pointwise relative error is not well defined in this context, since the current may vanish at certain spatial locations and time instants.
@@ 6.4.1 Within-the-distribution Validation
$\text{err}_{\lVert\cdot\rVert_{F}}$ $7.5194\cdot 10^{-2}$ $2.3042\cdot 10^{-2}$
@@ 6.4.2 Outside-the-distribution Validation
To evaluate the performance of the ROM s outside the training distribution, two additional transient simulations were performed at 50 Hz with excitation amplitudes of 10 mT and 30 mT–values lying below and above the training range (13–24 mT), respectively. The corresponding quantitative results, summarized in Table 3, indicate that the accuracy of the Structured Neural ODE decreases noticeably, whereas the POD-DEIM ROM maintains performance comparable to the within-distribution case.
@@
$\text{err}_{\lVert\cdot\rVert_{F}}$ $7.6155\cdot 10^{-2}$ $1.8361\cdot 10^{-1}$
@@ 6.4.3 Different Frequency Validation
To further assess the generalization capability of the proposed approach, additional simulations were carried out at an excitation amplitude of 20 mT–within the training range–but with frequencies of 40, 60, and 100 Hz, differing from the 50 Hz training frequency. The quantitative results, summarized in Table 4, demonstrate that the Structured Neural ODE achieves very good accuracy across all tested frequencies, with errors of the same order of magnitude as in the 50 Hz validation and consistently lower than those of the POD-DEIM ROM.
@@
$\text{err}_{\lVert\cdot\rVert_{F}}$ $7.5948\cdot 10^{-2}$ $3.6895\cdot 10^{-2}$
@@ 6.5 Computational Cost
To compare the computational cost of the two ROMs, we evaluate the number of Floating-Point Operations (FLOPs) required to advance the system by one time step. The time-stepping scheme used to simulate the POD-DEIM ROM employs Newton iterations at each step, with residual and Jacobian evaluations at each iteration–an approach necessitated by the strong nonlinearity of the problem. In both cases, an optimal implementation scenario is assumed.
@@ 6.5.2 Structure Neural ODE ROM Computational Cost
The computational cost for advancing the state in the Structured Neural ODE ROM includes evaluating the neural network, assembling the left- and right-hand sides through matrix-vector operations, and solving the resulting linear system.
@@ 6.5.3 Results of the Evaluations
For a single time step, the POD-DEIM ROM requires approximately $4.5\times 10^{7}$ operations, whereas the Structured Neural ODE ROM requires only $2.6\times 10^{5}$, corresponding to a nominal FLOPs-based speed-up of $\times 176$. This estimate reflects the algorithmic cost of advancing the reduced state under the fully implicit scheme employed in the POD-DEIM ROM, which requires repeated residual and Jacobian evaluations within Newton-Raphson iterations. In our experiments, the POD-DEIM ROM required an average of 19 Newton iterations per time-step, and the dominant cost arises from repeated assembly of the DEIM-reconstructed nonlinear term and its Jacobian. By contrast, the Structured Neural ODE ROM advances the reduced state by evaluating the neural-network once per step and solving a single reduced linear system, thereby avoiding Newton iterations and the associated repeated nonlinear assembly.
If, instead of considering operation counts, we measure computational cost in terms of wall-clock time for the full transient simulation, the observed speed is significantly larger. While the Structured Neural ODE can be implemented in a highly optimized manner using only efficient neural-network evaluation and basic algebraic operations, the POD-DEIM ROM requires repeated operator assembly, and these assembly routines are more difficult to optimize in practice. As a result, the overhead associated with repeated assembly and Newton iterations dominates the execution time of the POD-DEIM ROM. In our implementation, this leads to a practical speed-up of approximately $\times 1500$, reducing the runtime for a full transient simulation from 5 minutes to just 0.2 seconds. These results are summarized in Table 5.
@@
6.6 Time-step sensitivity analysis
We additionally investigate the effect of the time-step size $\Delta t$ on accuracy and computational cost. The Structured Neural ODE, trained at $\Delta t_{\text{train}}$, remains accurate when deployed with smaller time steps, whereas accuracy degrades for $\Delta t>\Delta t_{\text{train}}$, consistently with [13]. The fully implicit POD-DEIM scheme is, in principle, less sensitive to $\Delta t$, but in our experiments, Newton iterations fail to converge beyond a certain $\Delta t$, limiting the admissible time-step range unless additional solver enhancements are introduced (e.g., problem-specific preconditioning strategies and improved initial guesses), which bring non-trivial algorithmic complexity. These trends are summarized in Fig. 6.
Figure 6: Time-step sensitivity study. Each marker reports the mean absolute current error averaged over space and time with respect to the FOM reference solution, so that each point summarizes the discrepancy over the full transient. The dashed line denotes the training time step $\Delta t_{\text{train}}$. The DEIM curve terminates at the largest time step for which the Newton solver converges.
In terms of computational cost, the Structured Neural ODE exhibits an approximately constant per-step complexity (dominated by a single neural-network evaluation and reduced linear algebra), so the total transient evaluation time scales linearly with the number of time steps, i.e., proportional to $1/\Delta t$. The computational cost of the fully implicit POD-DEIM approach instead depends both on the number of time steps and on the number of Newton iterations per step. Moreover, as discussed in Section 6.7, the per-step cost of POD-DEIM with an explicit time integrator is already comparable to that of the Structured Neural ODE. Therefore, even in the most favorable implicit scenario (i.e., a single Newton iteration), one should expect a higher per-step cost due to the additional assembly of the Jacobian on top of the residual evaluation. Consequently, across the tested range of $\Delta t$, the proposed method retains a substantial computational advantage while maintaining competitive accuracy for $\Delta t\leq\Delta t_{\text{train}}$.
@@ 6.8 Post-processing: Power losses
As a post-processing step, we evaluate the AC losses in the CORC® cable, which serve as input for a thermal model aimed at predicting the cable’s temperature distribution–a key quantity for quench prediction in both design and control. The AC losses are computed as:
@@ 7 Discussion
Taken together, these results indicate that the Structured Neural ODE offers clear advantages in terms of intrusiveness and online efficiency, particularly in strongly nonlinear regimes where Newton iterations dominate the runtime. POD-DEIM remains preferable primarily in data-scarce or extrapolative scenarios, whereas the Structured Neural ODE is the method of choice whenever sufficient training data are available.